OpenStudy (anonymous):

In her first five basketball games, Tara scored 10, 11, 15, 16, and 18 points. What is the mean absolute deviation for this set of scores? points

OpenStudy (anonymous):

@StudyGurl14

OpenStudy (pinklion23):

i can help it is simple to find the mean

OpenStudy (ajanijones):

add them all, and divide by 5

OpenStudy (ajanijones):

Then you will find your answer. :)

OpenStudy (pinklion23):

yeah what he said or she idk

OpenStudy (pinklion23):

:/

OpenStudy (anonymous):

he and 14?

OpenStudy (studygurl14):

what @ajanijones said

OpenStudy (ajanijones):

, okay the mean is all the numbers added together and then divide by how many number there is. for your case: mean = 10+11+15+16+18=70 there is 5 numbers so... 70/5= 14 which you are correct. Hope this helps! I am a boy :)

OpenStudy (anonymous):

answer is 14?

OpenStudy (ajanijones):

Yep good job,

OpenStudy (pinklion23):

oh okay lol

OpenStudy (anonymous):

ty